# BRIDGE: Bootstrapping Text to Control Time-Series Generation via Multi-Agent Iterative Optimization and Diffusion Modeling

### v5

@@ 1 Introduction
(ii) Bridging discrepancy between text and time-series data for controlled TSG: Text and time-series (TS) data exhibit significant differences in both modality and granularity. Text conveys information via a fixed vocabulary of discrete tokens, while time series data is continuous, which leads to inherent mismatches. This disparity may render text too coarse to fully capture domain-specific patterns and characteristics, posing challenges in achieving precise domain-level control due to its incomplete or oversimplified representations. Meanwhile, text can provide detailed, instance-specific descriptions that are crucial for nuanced control, requiring careful and dedicated design to align text with TS features effectively.
@@ 4.2 Text Controlled Time Series Generation
Although text provides explicit details on trends, statistical properties, and domain-specific information, its discrete nature contrasts with the continuous structure of time-series data, posing challenges for using text as a control mechanism. To address this, we present our framework for text-controlled TSG. As shown in Figure 1, the framework employs a hybrid prompt that integrates semantic prototypes which supplement coarse-grained domain descriptions provided by text, along with textual descriptions. This design captures shared patterns across domains and improves the model’s generalization ability in unseen domains. We selected diffusion models (Ho et al., 2020; Lin et al., 2024) as the backbone of our framework due to their proven ability to generate diverse, high-quality data while effectively capturing complex data distributions. Recent successes of diffusion models in time-series forecasting (Rasul et al., 2021) and generation (Qian et al., 2024) further underscore their suitability for this task. The specific design of our architecture is detailed in Appendix B.2.
@@ 5.1.3 Evaluation Metrics
Assessing Controllability of TSG: To evaluate controllability in text-to-time series generation (TSG), we adopt the Joint Fréchet Time Series Distance (J-FTSD) (Narasimhan et al., 2024), which captures both local shape and global distributional similarity between the generated and reference sequences. This metric offers a more faithful measure of alignment than traditional point-wise metrics such as MSE or MAE, especially in cases where semantic similarity does not require exact numerical overlap. However, quantitative metrics alone may overlook perceptual or contextual alignment. To complement this, we conduct a Human Evaluation, where outputs are ranked based on how well they reflect the intended text descriptions. Annotators either evaluate samples independently (HE) or compare them among mixed distractors (HE@3), allowing for both absolute and relative quality assessment. This setup captures cases where generated outputs differ slightly (e.g., a constant shift) yet remain faithful in trend or pattern. Results are reported as average ranks across multiple evaluations. See Appendix H for details.
@@
Dataset Bridge Bridge w/o Text Bridge w/o Prototype TimeVQVAE TimeGAN GT-GAN TimeVAE Marginal Distribution Distance Electricity $0.220\pm 0.070$ $0.202\pm 0.066$ $0.277\pm 0.068$ $1.763\pm 0.088$ $2.443\pm 0.765$ $2.026\pm 0.280$ $3.306\pm 0.044$ Solar $375.531\pm 0.001$ $375.908\pm 10.230$ $376.111\pm 5.506$ $466.174\pm 0.145$ $460.810\pm 14.078$ $476.196\pm 17.041$ $365.906\pm 6.365$ Wind $0.316\pm 0.031$ $0.319\pm 0.046$ $0.362\pm 0.050$ $0.777\pm 0.028$ $1.115\pm 0.159$ $0.706\pm 0.106$ $0.943\pm 0.008$ Traffic $0.254\pm 0.034$ $0.261\pm 0.038$ $0.316\pm 0.006$ $1.170\pm 0.028$ $1.733\pm 0.137$ $1.311\pm 0.032$ $0.984\pm 0.012$ Taxi $0.386\pm 0.057$ $0.391\pm 0.057$ $0.491\pm 0.133$ $0.534\pm 0.032$ $1.278\pm 0.168$ $1.118\pm 0.157$ $0.697\pm 0.007$ Pedestrian $0.621\pm 0.124$ $0.624\pm 0.116$ $0.800\pm 0.111$ $1.225\pm 0.060$ $1.574\pm 0.290$ $1.559\pm 0.117$ $0.777\pm 0.012$ Air $0.447\pm 0.112$ $0.410\pm 0.129$ $0.508\pm 0.091$ $0.338\pm 0.012$ $2.089\pm 0.618$ $2.828\pm 0.172$ $1.369\pm 0.040$ Temperature $0.342\pm 0.010$ $0.345\pm 0.019$ $0.408\pm 0.053$ $0.943\pm 0.035$ $1.164\pm 0.110$ $1.165\pm 0.072$ $2.044\pm 0.024$ Rain $5.340\pm 0.421$ $5.597\pm 0.409$ $6.678\pm 2.045$ $9.243\pm 0.122$ $10.937\pm 4.039$ $6.473\pm 1.207$ $9.134\pm 0.477$ NN5 $0.591\pm 0.029$ $0.628\pm 0.021$ $0.748\pm 0.134$ $1.424\pm 0.043$ $2.758\pm 0.142$ $2.121\pm 0.094$ $2.871\pm 0.045$ Fred-MD $0.258\pm 0.045$ $0.271\pm 0.056$ $0.359\pm 0.097$ $2.932\pm 0.133$ $4.028\pm 0.130$ $4.026\pm 0.087$ $2.902\pm 0.215$ Exchange $0.374\pm 0.053$ $0.376\pm 0.058$ $0.382\pm 0.041$ $0.993\pm 0.058$ $1.553\pm 0.122$ $1.355\pm 0.072$ $1.331\pm 0.042$ K-L Divergence Electricity $0.011\pm 0.010$ $0.014\pm 0.013$ $0.013\pm 0.005$ $0.185\pm 0.018$ $0.395\pm 0.121$ $0.415\pm 0.040$ $0.580\pm 0.005$ Solar $0.007\pm 0.002$ $0.007\pm 0.003$ $0.015\pm 0.012$ $0.726\pm 0.043$ $0.889\pm 0.288$ $0.102\pm 0.045$ $0.201\pm 0.008$ Wind $0.067\pm 0.030$ $0.061\pm 0.042$ $0.069\pm 0.030$ $0.493\pm 0.081$ $4.528\pm 1.743$ $0.511\pm 0.129$ $0.553\pm 0.014$ Traffic $0.013\pm 0.004$ $0.013\pm 0.005$ $0.017\pm 0.004$ $0.145\pm 0.015$ $2.134\pm 0.952$ $1.108\pm 0.171$ $0.212\pm 0.006$ Taxi $0.013\pm 0.009$ $0.013\pm 0.010$ $0.039\pm 0.031$ $0.100\pm 0.014$ $1.160\pm 0.651$ $0.663\pm 0.127$ $0.120\pm 0.005$ Pedestrian $0.011\pm 0.009$ $0.011\pm 0.009$ $0.020\pm 0.005$ $0.275\pm 0.021$ $0.881\pm 0.436$ $0.347\pm 0.085$ $0.052\pm 0.010$ Air $0.022\pm 0.017$ $0.018\pm 0.014$ $0.019\pm 0.011$ $0.017\pm 0.004$ $0.588\pm 0.369$ $0.506\pm 0.091$ $0.176\pm 0.016$ Temperature $0.023\pm 0.013$ $0.024\pm 0.010$ $0.039\pm 0.033$ $0.980\pm 0.190$ $8.775\pm 2.511$ $2.177\pm 0.323$ $1.910\pm 0.076$ Rain $0.006\pm 0.001$ $0.007\pm 0.001$ $0.009\pm 0.003$ $0.008\pm 0.002$ $0.383\pm 0.089$ $0.462\pm 0.056$ $0.175\pm 0.011$ NN5 $0.010\pm 0.008$ $0.013\pm 0.008$ $0.016\pm 0.007$ $0.603\pm 0.107$ $4.054\pm 1.592$ $1.372\pm 0.180$ $1.284\pm 0.058$ Fred-MD $0.024\pm 0.019$ $0.028\pm 0.029$ $0.033\pm 0.029$ $0.712\pm 0.054$ $5.371\pm 1.455$ $3.509\pm 0.299$ $0.376\pm 0.025$ Exchange $0.083\pm 0.056$ $0.088\pm 0.065$ $0.088\pm 0.077$ $1.984\pm 0.836$ $4.376\pm 0.664$ $1.583\pm 0.932$ $2.011\pm 0.433$
@@ 5.3.1 Performance on In-domain settings
Electricity Solar Wind Traffic J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 Bridge $0.538$ $2.3$ $2.4$ $0.295$ $2.6$ $2.8$ $5.011$ $1.8$ $1.8$ $0.570$ $1.4$ $1.4$ w/o Prototype $1.164$ $3.3$ $3.4$ $0.322$ $3.2$ $3.3$ $6.843$ $3.1$ $3.2$ $0.597$ $2.1$ $2.1$ w/o Text $1.821$ $3.4$ $3.5$ $0.330$ $3.2$ $3.4$ $6.935$ $4.1$ $4.2$ $0.611$ $4.2$ $4.4$ Taxi Pedestrian Air Temperature J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 Bridge $0.974$ $1.4$ $1.4$ $0.488$ $2.2$ $2.6$ $0.654$ $2.7$ $2.8$ $3.977$ $2.4$ $2.4$ w/o Prototype $1.037$ $3.4$ $3.6$ $0.662$ $2.6$ $3.1$ $0.817$ $3.0$ $3.1$ $4.613$ $2.6$ $3.1$ w/o Text $1.312$ $4.2$ $4.4$ $0.550$ $3.3$ $5.0$ $0.677$ $3.3$ $3.4$ $4.708$ $3.6$ $3.9$ Rain NN5 Fred-MD Exchange J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 J-FTSD HE HE@3 Bridge $0.147$ $1.9$ $1.9$ $0.972$ $2.5$ $2.8$ $0.260$ $2.5$ $2.8$ $1.581$ $1.9$ $1.9$ w/o Prototype $0.141$ $3.2$ $3.2$ $1.115$ $3.1$ $3.5$ $0.341$ $2.7$ $3.0$ $1.846$ $3.1$ $3.1$ w/o Text $0.151$ $3.7$ $3.7$ $1.192$ $3.4$ $3.8$ $0.423$ $3.8$ $4.2$ $1.738$ $3.7$ $3.7$
@@
As shown in Table 3, the proposed Bridge consistently outperforms the two ablated variants in most of datasets, according to both the J-FTSD score and human evaluation metrics (HE and HE@3). The removal of textual input leads to the most severe performance drop, particularly in human assessments—demonstrating that text is essential for producing outputs that align with intended semantics. Without textual guidance, the model tends to generate less coherent and less interpretable sequences. For example, on the Traffic and Pedestrian datasets, HE scores increase by over 3 points without text, indicating poor alignment. The “w/o Prototype” setting shows moderate degradation, suggesting that while prototype representations enhance fine-grained alignment and structure, they are less critical than textual descriptions. Notably, the gap is especially visible on datasets with high temporal variability such as Wind and Temperature. These results confirm that both components—text and prototypes—play complementary roles in generating realistic and controllable time series.
@@ 5.3.2 Performance on Unseen domain settings
Table 4 highlights the structural contributions of text and prototype components in unseen domains. The full Bridge model achieves the lowest J-FTSD and most favorable human evaluation scores across both Stock and Web datasets, indicating strong generalization beyond the training distribution. Removing text input causes the most notable performance drop, especially in J-FTSD and HE@3, confirming that textual guidance is critical for semantic alignment and coherent generation. In comparison, the “w/o Prototype” variant shows only moderate degradation, particularly in HE but not in J-FTSD, suggesting that prototypes help enforce structural consistency but are less essential than text for global control. These results support a complementary division of labor: text governs high-level semantics, while prototypes refine local dynamics.
Stock Web J-FTSD HE HE@3 J-FTSD HE HE@3 Bridge $7.483$ $2.8$ $2.8$ $5.529$ $2.7$ $3.2$ w/o Prototype $7.687$ $2.8$ $3.5$ $5.752$ $3.1$ $3.5$ w/o Text $8.178$ $3.4$ $4.0$ $6.302$ $3.2$ $3.4$
@@ 5.4 Performance Analysis on Controlling Text
Text Types AirPassenger Sunspots LLMTime LSTPrompt LLMTime LSTPrompt Rule-based Text $52.41$ $20.08$ $63.92$ $51.61$ Initial Text $49.36$ $15.12$ $59.88$ $49.71$ Refined Text(Ours) $40.94$ $12.39$ $48.64$ $42.37$hdashlinew/o Instance Context $41.96$ $13.54$ $54.33$ $44.23$ w/o Background $44.63$ $14.77$ $56.81$ $46.07$ w/o Statistical Context $44.01$ $13.41$ $54.24$ $47.12$ w/o Pattern $44.36$ $14.52$ $55.16$ $46.84$ w/o Pattern+Statistic $44.30$ $14.27$ $56.89$ $45.65$hdashlineBaseline (Liu et al., 2024b) $45.75$ $15.00$ $59.91$ $47.59$

### v6

@@ 1 Introduction
(ii) Bridging discrepancy between text and time-series data for controlled TSG:
Text and time-series (TS) data exhibit significant differences in both modality and granularity. Text conveys information via a fixed vocabulary of discrete tokens, while time series data is continuous, which leads to inherent mismatches. This disparity may render text too coarse to fully capture domain-specific patterns and characteristics, posing challenges in achieving precise domain-level control due to its incomplete or oversimplified representations. Meanwhile, text can provide detailed, instance-specific descriptions that are crucial for nuanced control, requiring careful and dedicated design to align text with TS features effectively.
@@ 4.2 Text Controlled Time Series Generation
Although text provides explicit details on trends, statistical properties, and domain-specific information, its discrete nature contrasts with the continuous structure of time-series data, posing challenges for using text as a control mechanism. To address this, we present our framework for text-controlled TSG. As shown in Figure 1, the framework employs a hybrid prompt that integrates semantic prototypes which supplement coarse-grained domain descriptions provided by text, along with textual descriptions. This design captures shared patterns across domains and improves the model’s generalization ability in unseen domains.
We selected diffusion models (Ho et al., 2020; Lin et al., 2024) as the backbone of our framework due to their proven ability to generate diverse, high-quality data while effectively capturing complex data distributions. Recent successes of diffusion models in time-series forecasting (Rasul et al., 2021) and generation (Qian et al., 2024) further underscore their suitability for this task. The specific design of our architecture is detailed in Appendix B.2.
@@ 5.1.3 Evaluation Metrics
Assessing Controllability of TSG: To evaluate controllability in text-to-time series generation (TSG), we adopt the Joint Fréchet Time Series Distance (J-FTSD) (Narasimhan et al., 2024), which captures both local shape and global distributional similarity between the generated and reference sequences. This metric offers a more faithful measure of alignment than traditional point-wise metrics such as MSE or MAE, especially in cases where semantic similarity does not require exact numerical overlap. However, quantitative metrics alone may overlook perceptual or contextual alignment. To complement this, we conduct a Human Evaluation, where outputs are ranked based on how well they reflect the intended text descriptions. We consider two settings: in HE-Rank, annotators evaluate only the outputs generated by different model settings along with the ground truth time series; in HE-Mixed, these candidates are presented alongside three randomly selected distractor time series from the test set. This design enables both focused and distractor-aware evaluation, capturing subtle differences (e.g., constant shifts) while assessing trend or pattern alignment. Results are reported as average ranks across multiple evaluations. See Appendix H for details.
@@
Dataset Bridge Bridge w/o Text Bridge w/o Prototype TimeVQVAE TimeGAN GT-GAN TimeVAE Marginal Distribution Distance Electricity $20\pm 0.070 0.202$$02\pm 0.066$ $0.277\pm 0.068$ $1.763\pm 0.088$ $2.443\pm 0.765$ $2.026\pm 0.280$ $3.306\pm 0.044$ Solar $375.531\pm 0.001$ $375.908\pm 10.230$ $376.111\pm 5.506$ $466.174\pm 0.145$ $460.810\pm 14.078$ $476.196\pm 17.041$ $365.906\pm 6.365$ Wind $0.316\pm 0.031$ $0.319\pm 0.046$ $0.362\pm 0.050$ $0.777\pm 0.028$ $1.115\pm 0.159$ $0.706\pm 0.106$ $0.943\pm 0.008$ Traffic $0.254\pm 0.034$ $0.261\pm 0.038$ $0.316\pm 0.006$ $1.170\pm 0.028$ $1.733\pm 0.137$ $1.311\pm 0.032$ $0.984\pm 0.012$ Taxi $0.386\pm 0.057$ $0.391\pm 0.057$ $0.491\pm 0.133$ $0.534\pm 0.032$ $278\pm 0.168 1.1$$118\pm 0.157$ $0.697\pm 0.007$ Pedestrian $0.621\pm 0.124$ $0.624\pm 0.116$ $0.800\pm 0.111$ $1.225\pm 0.060$ $1.574\pm 0.290$ $1.559\pm 0.117$ $0.777\pm 0.012$ Air $0.447\pm 0.112$ $0.410\pm 0.129$ $0.508\pm 0.091$ $0.338\pm 0.012$ $2.089\pm 0.618$ $2.828\pm 0.172$ $1.369\pm 0.040$ Temperature $0.342\pm 0.010$ $0.345\pm 0.019$ $0.408\pm 0.053$ $0.943\pm 0.035$ $164\pm 0.110 1.1$$165\pm 0.072$ $2.044\pm 0.024$ Rain $340\pm 0.421 5.5$$597\pm 0.409$ $6.678\pm 2.045$ $9.243\pm 0.122$ $10.937\pm 4.039$ $6.473\pm 1.207$ $9.134\pm 0.477$ NN5 $0.591\pm 0.029$ $0.628\pm 0.021$ $0.748\pm 0.134$ $1.424\pm 0.043$ $2.758\pm 0.142$ $2.121\pm 0.094$ $2.871\pm 0.045$ Fred-MD $0.258\pm 0.045$ $0.271\pm 0.056$ $0.359\pm 0.097$ $2.932\pm 0.133$ $4.028\pm 0.130$ $4.026\pm 0.087$ $2.902\pm 0.215$ Exchange $0.374\pm 0.053$ $0.376\pm 0.058$ $0.382\pm 0.041$ $0.993\pm 0.058$ $1.553\pm 0.122$ $55\pm 0.072 1.331$$31\pm 0.042$ K-L Divergence Electricity $011\pm 0.010 0.0$$14\pm 0.013 0.0$$013\pm 0.005$ $0.185\pm 0.018$ $0.395\pm 0.121$ $0.415\pm 0.040$ $0.580\pm 0.005$ Solar $007\pm 0.002 0.0$$07\pm 0.003 0.0$$015\pm 0.012$ $0.726\pm 0.043$ $0.889\pm 0.288$ $0.102\pm 0.045$ $0.201\pm 0.008$ Wind $067\pm 0.030 0.0$$61\pm 0.042 0.0$$069\pm 0.030$ $0.493\pm 0.081$ $4.528\pm 1.743$ $0.511\pm 0.129$ $0.553\pm 0.014$ Traffic $013\pm 0.004 0.0$$13\pm 0.005 0.0$$017\pm 0.004$ $0.145\pm 0.015$ $2.134\pm 0.952$ $1.108\pm 0.171$ $0.212\pm 0.006$ Taxi $013\pm 0.009 0.0$$13\pm 0.010 0.0$$039\pm 0.031$ $0.100\pm 0.014$ $1.160\pm 0.651$ $0.663\pm 0.127$ $0.120\pm 0.005$ Pedestrian $1\pm 0.009 0.011±0.0$$1\pm 0.009$ $0.020\pm 0.005$ $0.275\pm 0.021$ $0.881\pm 0.436$ $347\pm 0.085 0.0$$052\pm 0.010$ Air $022\pm 0.017 0.0$$18\pm 0.014 0.0$$19\pm 0.011 0.0$$017\pm 0.004$ $0.588\pm 0.369$ $0.506\pm 0.091$ $0.176\pm 0.016$ Temperature $023\pm 0.013 0.0$$24\pm 0.010 0.0$$039\pm 0.033$ $0.980\pm 0.190$ $8.775\pm 2.511$ $2.177\pm 0.323$ $1.910\pm 0.076$ Rain $006\pm 0.001 0.0$$07\pm 0.001 0.0$$09\pm 0.003 0.0$$008\pm 0.002$ $0.383\pm 0.089$ $0.462\pm 0.056$ $0.175\pm 0.011$ NN5 $010\pm 0.008 0.0$$13\pm 0.008 0.0$$016\pm 0.007$ $0.603\pm 0.107$ $4.054\pm 1.592$ $1.372\pm 0.180$ $1.284\pm 0.058$ Fred-MD $024\pm 0.019 0.0$$28\pm 0.029 0.0$$033\pm 0.029$ $0.712\pm 0.054$ $5.371\pm 1.455$ $3.509\pm 0.299$ $0.376\pm 0.025$ Exchange $3\pm 0.056 0.088±0.0$$8\pm 0.065$ $0.088\pm 0.077$ $1.984\pm 0.836$ $4.376\pm 0.664$ $1.583\pm 0.932$ $2.011\pm 0.433$
@@ 5.3.1 Performance on In-domain settings
Electricity Solar Wind Traffic J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed Bridge $0.538$ $2.3$ $2.4$ $0.295$ $2.6$ $2.8$ $5.011$ $1.8$ $1.8$ $0.570$ $1.4$ $1.4$ w/o Prototype $1.164$ $3.3$ $3.4$ $0.322$ $3.2$ $3.3$ $6.843$ $3.1$ $3.2$ $0.597$ $2.1$ $2.1$ w/o Text $1.821$ $3.4$ $3.5$ $0.330$ $3.2$ $3.4$ $6.935$ $4.0$ $4.2$ $0.611$ $4.2$ $4.4$ Taxi Pedestrian Air Temperature J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed Bridge $0.974$ $1.4$ $1.4$ $0.488$ $2.2$ $2.6$ $0.654$ $2.7$ $2.8$ $3.977$ $2.4$ $2.4$ w/o Prototype $1.037$ $3.4$ $3.6$ $0.662$ $2.6$ $3.1$ $0.817$ $3.0$ $3.1$ $4.613$ $2.6$ $3.1$ w/o Text $1.312$ $4.2$ $4.4$ $0.550$ $3.3$ $4.0$ $0.677$ $3.3$ $3.4$ $4.708$ $3.6$ $3.9$ Rain NN5 Fred-MD Exchange J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed Bridge $0.147$ $1.9$ $1.9$ $0.972$ $2.5$ $2.8$ $0.260$ $2.5$ $2.8$ $1.581$ $1.9$ $1.9$ w/o Prototype $0.141$ $3.2$ $3.2$ $1.115$ $3.1$ $3.5$ $0.341$ $2.7$ $3.0$ $1.846$ $3.1$ $3.1$ w/o Text $0.151$ $3.7$ $3.7$ $1.192$ $3.4$ $3.8$ $0.423$ $3.8$ $4.2$ $1.738$ $3.7$ $3.7$
@@
As shown in Table 3, the proposed Bridge consistently outperforms the two ablated variants in most of datasets, according to both the J-FTSD score and human evaluation metrics (HE and HE+3). The removal of textual input leads to the most severe performance drop, particularly in human assessments—demonstrating that text is essential for producing outputs that align with intended semantics. Without textual guidance, the model tends to generate less coherent and less interpretable sequences. For example, on the Traffic and Pedestrian datasets, HE scores increase by over 3 points without text, indicating poor alignment. The “w/o Prototype” setting shows moderate degradation, suggesting that while prototype representations enhance fine-grained alignment and structure, they are less critical than textual descriptions. Notably, the gap is especially visible on datasets with high temporal variability such as Wind and Temperature. These results confirm that both components—text and prototypes—play complementary roles in generating realistic and controllable time series.
@@ 5.3.2 Performance on Unseen domain settings
Table 4 highlights the structural contributions of text and prototype components in unseen domains. The full Bridge model achieves the lowest J-FTSD and most favorable human evaluation scores across both Stock and Web datasets, indicating strong generalization beyond the training distribution. Removing text input causes the most notable performance drop, especially in J-FTSD and HE+3, confirming that textual guidance is critical for semantic alignment and coherent generation. In comparison, the “w/o Prototype” variant shows only moderate degradation, particularly in HE but not in J-FTSD, suggesting that prototypes help enforce structural consistency but are less essential than text for global control. These results support a complementary division of labor: text governs high-level semantics, while prototypes refine local dynamics.
Stock Web J-FTSD HE-Rank HE-Mixed J-FTSD HE-Rank HE-Mixed Bridge $7.483$ $2.8$ $2.8$ $5.529$ $2.7$ $3.2$ w/o Prototype $7.687$ $2.8$ $3.5$ $5.752$ $3.1$ $3.5$ w/o Text $8.178$ $3.4$ $4.0$ $6.302$ $3.2$ $3.4$
@@ 5.4 Performance Analysis on Controlling Text
Text Types AirPassenger Sunspots LLMTime LSTPrompt LLMTime LSTPrompt Rule-based Text $52.41$ $20.08$ $63.92$ $51.61$ Initial Text $49.36$ $15.12$ $59.88$ $49.71$ Refined Text(Ours) $40.94$ $12.39$ $48.64$ $42.37$ w/o Instance Context $41.96$ $13.54$ $54.33$ $44.23$ w/o Background $44.63$ $14.77$ $56.81$ $46.07$ w/o Statistical Context $44.01$ $13.41$ $54.24$ $47.12$ w/o Pattern $44.36$ $14.52$ $55.16$ $46.84$ w/o Pattern+Statistic $44.30$ $14.27$ $56.89$ $45.65$ Baseline (Liu et al., 2024b) $45.75$ $15.00$ $59.91$ $47.59$